Please answer these questions, accurately showing all steps. Thanks. Ensure you are giving the correct answer.

Problem B Seattle Cycle, founded in 2016, manufactures a racing bike. In 2016 and 2017, Bart Wilson was the general manager. The owner was very pleased with the financial performance of the company during those years. The company had a loss in the first year of operation but had a small profit in 2017. Bart quit at the end of 2017. His replacement, Chad Nguyen, did not appear to be as successful. Chad argued that he was improving the company by getting rid of excess inventory and increasing sales volume. The owner only saw the big loss during Chads first year as general manager 2016 2017 2018 Production (units) Sales (units) 1,000 1,150000 Sales price Variable product cost per unit Total cost per unit (absorption) Total sales & admin costs (all fixed) 1,000 1,000 1,000 400400 400 $880 800 1,000 $200,000 200,000 200,000 Full absorption method income statements: 2016 2017 2018 Total Sales COGS $1,000,000 $1,150,000 $1,500,000 3,650,000 Beginning inventory 220,000 480,000 1,100,000 1,200,000 1,000,000 3,300,000 700,000 COG manufactured Less: Ending inventory Gross margin Selling & administrative exp (220,000 880,000 940,000 1,380,000 3,200,000 120,000 210,000 (100 120,000 200,000 480 (800 200,000 200,000 10,000 600,000 (150,000) Net income (loss) (80,000) Assume Seattle Cycle had used variable costing during the period. What would NOI be for each of the three years under variable costing? 2016 2017 2018 13 How much FMOH did Seattle Cycle defer in ending inventory in the following years? 2016 2017 14 16 After looking at your results, what advice might you give the owner of Seattle Cycle? (Assume that the owner has only seen the full absorption income statements.) a. Bart did a great job of building inventory, resulting in better operating results. Rehire him. b. Although it looks like Bart generated better results when full absorption is used, Chads focus on building sales and controlling inventory will pay off in the long run. c. Chads focus on building sales and controlling inventory sounds good but if it doesnt result in higher net income under GAAP, the focus needs to change.
Answer #1

Q1) The Net Operating Income(NOI)

for the year 2016:

  • (Sales units X Sales Price) = 10,00,000 $
  • 250 Unit Inventory
  • (Cost per unit) X (Production Unit) = 11,00,000 $
  • 10,00,000 $ - 11,00,000 $ = - 1,00,000 $ (Profit)

for the year 2017:

  • (Sales units X Sales Price) = 11,50,000 $
  • 350 Unit Inventory
  • (Cost per unit) X (Production Unit) = 12,00,000 $
  • 11,50,000 $ - 12,00,000 $= - 50,000 $ (Profit)

for the year 2018:

  • (Sales units X Sales Price) = 15,00,000 $
  • Previous Years 250+350 Unit Inventory = 600 (500 sold in the current year)
  • (Cost per unit) X (Production Unit) = 10,00,000 $
  • 15,00,000 $ - 10,00,000 $= 5,00,000 $ (Profit)

Q2) Fixed Manufacturing Overhead

for the year 2016:

  • (Variable Product cost) - (Total Product Cost) = 480 $ X 1500 (Unit) = 6,00,000 $
  • (Total Cost) + (Inventory) = 14,80,000 $

for the year 2017:

  • (Variable Product cost) - (Total Product Cost) = 400 $ X 1500 (Unit) = 6,00,000 $
  • (Total Cost) + (Inventory) = 15,40,000 $

Q3) Although it looks like Bart generated better result when full absorption is used, chad's focus on building sales and controlling inventory will payoff in the long run.

As per 2016 reports Bart and incurred lot of cost by producing a lot of goods which went into inventory as batch manufacturing and selling at a higher margin. As per 2017 report Chad started slowly clearing inventory and increase the volume of sales. As inventory incurs cost of storage and chance of damage of goods, maintaining a parity between sales at a lesser margin and producing what you can sell helps in generating more profit for a business in the long run.